OpenStudy (anonymous):

Jack bought 3 slices of cheese pizza and 4 slices of mushroom pizza for a total cost of $12.50. Grace bought 3 slices of cheese pizza and 2 slices of mushroom pizza for a total cost of $8.50. What is the cost of one slice of mushroom pizza? a) $1.50 b) $2.00 c) $3.00 d) $3.50

OpenStudy (amistre64):

3c + 4m = 12.50 3c + 2m = 8.50 3c + 4m = 12.50 -3c -2m = -8.50 ------------------ 2m = 4.00 m = 2.00